### v1

# Probing the band structure of the strongly correlated antiferromagnet NiPS3 across its phase transition

Abstract
NiPS3 is an exfoliable van-der-Waals intralayer antiferromagnet with zigzag-type spin arrangement. It is distinct from other TMPS3 (TM: transition metal) materials by optical excitations into a strongly correlated state that is tied to the magnetic properties. However, the related, fundamental band structure across the antiferromagnetic phase transition has not been probed yet. Here, we use angular-resolved photoelectron spectroscopy with m resolution in combination with DFT+U calculations for that purpose. We identify a characteristic band shift across . It is attributed to bands of mixed Ni and S character related to the superexchange interaction of Ni 3t2g orbitals. Moreover, we find a structure above the valence band maximum with little angular dispersion that could not be reproduced by the calculations. The discrepancy suggests the influence of many-body interactions beyond the DFT+U approximations in striking contrast to the results on MnPS3 and FePS3, where these calculations were sufficient for an adequate description.
†These authors contributed equally to this work. ∗Corresponding author: M. Morgenstern, Email: mmorgens@physik.rwth-aachen.de
I Introduction
Van der Waals (vdW)-type two-dimensional (2D) ferromagnets (FMs) and antiferromagnets (AFMs) Huang et al. (2017); Sun et al. (2019); Gong et al. (2017); Wang et al. (2018a); Deng et al. (2018); Fei et al. (2018) enrich the possibilities for spintronics, magneto-optical devices and magnetic sensors Li et al. (2019); Zhang et al. (2019); Burch et al. (2018); Gong and Zhang (2019); Huang et al. (2020). This is related to the fact that the 2D magnetism can be tuned via magnetic fields, electrostatic gating, current pulses, strain, light, ion intercalation, proximity and moiré lattices Burch et al. (2018); Li et al. (2019); Xu et al. (2021); Mi et al. (2022); Shin et al. (2022); Ahn et al. (2024); Cham et al. (2024) as well as due to the relatively strong magneto-electric, magneto-elastic and magneto-optic couplings Wang et al. (2022); Vaclavkova et al. (2021); Ressouche et al. (2010); Kang et al. (2020); Hwangbo et al. (2021); Dirnberger et al. (2022). Moreover, these materials enable a novel access to fundamental questions in 2D magnetism such as the Berezinskii–Kosterlitz–Thouless (BKT) transition or proximity-induced modifications of the magnetization Gong et al. (2017); Burch et al. (2018); Gong and Zhang (2019); Huang et al. (2020); Srivastava et al. (2020); Bedoya-Pinto et al. (2021); Albarakati et al. (2022); Ahn et al. (2024); Sun et al. (2024).
A unique type are transition metal (TM) phosphorus trisulfides TMPS3 Joy and Vasudevan (1992); Wang et al. (2018b). In contrast to other vdW magnets, they are intralayer AFMs and, moreover, rather directly realize the three most prominent classes of 2D magnetism (Ising, XY, Heisenberg) Chittari et al. (2016); Kim et al. (2019); Autieri et al. (2022). Hence, the materials provide a versatile platform to study 2D AFM properties. The TMPS3 materials are semiconductors with a layered honeycomb structure of the TM2+ atoms, each surrounded by covalently bonded (P2S6)4- bipyramids (Fig. 1a). The magnetism arises from the competition between direct exchange interactions between neighboring TM atoms and indirect superexchange mediated by S and P atoms Autieri et al. (2022); Yan et al. (2023a); Rybak et al. (2024). Depending on the relative strength of the competitive interactions, various long-range AFM orders have been found, dubbed zigzag, Néel, and stripy AFM order Joy and Vasudevan (1992); Sivadas et al. (2015).
NiPS3 stands out since it exhibits significant spin-charge coupling, leading to spin-correlated excitons whose origin is actively debated. The excitons display layer-dependent characteristics and polarization properties correlated with the material’s magnetism Kim et al. (2018); Kang et al. (2020); Lane and Zhu (2020); Wang et al. (2021); Hwangbo et al. (2021); Wang et al. (2022); Jana et al. (2023); Klaproth et al. (2023); Song et al. (2024); Wang et al. (2024); Allington et al. (2025). Notably, an ultra-sharp photoluminescence peak has been observed at approximately 1.47 eV Chittari et al. (2016); Kim et al. (2018); Muhammad et al. (2023); Du et al. (2016). Moreover, the magnetic interactions are surprisingly dominated by a third nearest neighbor AFM exchange coupling due to a substantial overlap of p- and d-orbitals along the corresponding Ni-S-P-S-Ni interaction path Scheie et al. (2023); Autieri et al. (2022); Chittari et al. (2016); Yan et al. (2023a). This leads to an intralayer AFM zigzag order as the ground state below the bulk Néel temperature 155 K, where the spins are oriented mostly in-plane along the zigzag direction (Fig. 1a) with a slight canting out-of-plane by 8∘ Wang et al. (2022). In monolayer form, the long-range magnetic order is most likely suppressed potentially leading to a BKT transition Kim et al. (2019); Sun et al. (2024). This behavior aligns with descriptions based on a highly anisotropic XXZ model Joy and Vasudevan (1992); Wildes et al. (2015).
The magnetic ordering of NiPS3 has already been probed by multiple methods Kim et al. (2018, 2019); Kang et al. (2020); Hwangbo et al. (2021); Afanasiev et al. (2021); Basnet et al. (2021, 2022); Scheie et al. (2023); Muhammad et al. (2023), but none of them is sensitive to details of the band structure, which has only been probed at room temperature so far, i. e. above , using angular resolved photoelectron spectroscopy (ARPES) Yan et al. (2023b); Klaproth et al. (2023); Nitschke et al. (2025). Here, we provide band structure mapping by ARPES at variable temperature above and below leveraging the approach that we recently developed for the isostructural MnPS3 and FePS3 Strasdas et al. (2023); Pestka et al. (2024). We prevent sample charging due to photoelectrons, that often appears for semiconductors at low Bianchi et al. (2023); De Vita et al. (2022); Nitschke et al. (2023); Koitzsch et al. (2023); Voloshina et al. (2023), by exfoliating thin films of NiPS3 onto a conductive Au film deposited on Si/SiO2, itself necessary for identifying the flakes optically. The resulting band structure is compared with density functional theory calculations (DFT+U) enabling the identification of the elemental and orbital character of the different bands and revealing an optimal Hubbard parameter eV. Most importantly, we observe a pronounced band shift across for a valence band identified as a mixed Ni 3d/S 3p band that could be attributed to a next nearest neighbor superexchange path. Moreover, we reveal an additional band feature in ARPES that appears above the valence band maximum, but could not be reproduced by the DFT+U calculations. We believe that it testifies the strong correlations in NiPS3 compared to MnPS3 and FePS3 in line with the optical results Lane and Zhu (2020); Wang et al. (2022); Kim et al. (2018); Kang et al. (2020); Hwangbo et al. (2021).
II Results and Discussion
NiPS3 single crystals are synthesized by the vapor transport method. Flakes are subsequently exfoliated by the scotch tape method onto Au/Ti covered Si/SiO2 at 60∘ C, shortly after the substrate was cleaned by plasma ashing Strasdas et al. (2023) (Materials and Methods). Flakes with appropriate thickness and lateral size are identified using optical and atomic force microscopy (Fig. 1b-c). Typically, these flakes exhibit various thicknesses being most thin at the rim. In this paper, we studied a 15 layer thick area with a lateral size of m2 as marked in Fig. 1b. Atomic force microscopy reveals a rms roughness of nm and no obvious contaminations in that area (Fig. 1c). The relatively small thickness is crucial to avoid charging during -ARPES at low temperature as detected, e. g. for 45 layers at 90 K, where all bands are shifted downwards by 0.3 eV.
Figure 1d displays X-ray photoelectron spectroscopy (XPS) data of the 15 layer area. All components of NiPS3 appear as core level peaks Yan et al. (2021); Currò et al. (1995); Piacentini et al. (1982) as well as a S Auger peak Coad and Rivière (1972). The latter shifts with respect to the other peaks by changing the photon energy exposing the Ni 3p peak Justin Gorham (2012). A small contribution from the spin-split Au 4f peaks Justin Gorham (2012) is also visible likely originating from the substrate due to the relatively large spot of the photon beam on the sample (diameter: 15 m) at the large for XPS.
The antiferromagnetic phase transition of the exfoliated NiPS3 flakes is verified by Raman spectroscopy (Fig. 1e). Due to the nearly realized D3d intralayer point group symmetry Kuo et al. (2016), NiPS3 has eight major Raman-active phonon modes, all visible in our Raman spectra with frequencies in good agreement with the literature Kim et al. (2019); Wang et al. (2019); Muhammad et al. (2023); Kuo et al. (2016). We reproduce three signatures that have previously been assigned to the AFM phase transition Kim et al. (2019); Muhammad et al. (2023). Firstly, a broad 2-magnon scattering peak, centered at 530 cm-1, appears below at 77 K and is absent at room temperature (RT) Kim et al. (2019). Secondly, the peak P2 caused by in-plane vibrations of the Ni atoms splits into a double peak (insets, Fig. 1e) as attributed to a pronounced magnon-phonon coupling Kim et al. (2019); Muhammad et al. (2023). Thirdly, the peak P8, caused by an in-plane Eg-type vibration of the (P2S6)4- bipyramids, develops a Fano-resonance line shape below T due to an interference between the phonon excitation and the continuum of 2-magnon excitations Fano (1961). These three fingerprints confirm the AFM phase transition in our NiPS3 flakes.
Figure 2 shows the main result of this work, namely a band shift across T near the point. We present the raw ARPES data (a–b), the data after box-type smoothing (c, only the energy distribution curve (EDC) is shown), after additional background subtraction (d–f) and angularly averaged in -space around the point (g–h). Already the raw data (Fig. 2a–b) exhibit a slightly enhanced intensity below in between the white arrows. This change can be identified as an additional peak in the EDC at after mild smoothing (upper arrow, Fig. 2c) implying a state at eV below (: energy, : Fermi energy as probed on Au). A polynomial background subtraction (Fig. 2f) increases the visibility of that peak. The resulting energy vs. wave vector display of the photoelectron intensity (Fig. 2d–e) reveals that the peak belongs to a rather flat band extending up to an in-plane wave vector Å, where it merges with upwards dispersing bands. These features are corroborated by the angulary averaged data in space (Fig. 2g–h). The band at eV is again visible below (pink arrow, Fig. 2h). It shifts downwards by roughly 150 meV above (Fig. 2g), nearly merging with the band at eV. The insets finally show the two-dimensional plots at that energy exhibiting a ring above due to the upwards dispersing bands and a full circle below due to the additional flat band. The upwards dispersing bands visible in the main image are also slightly less steep below .
Next, we compare the ARPES data with DFT+U calculations of the magnetic ground state (AFM zigzag type) and other magnetic configurations (AFM Néel and disordered). This aims to identify the elemental and orbital character of the different bands, in particular the ones that are changing across . We started by selecting the proper Hubbard parameter (DFT+U) by detailed comparison with the experimental data. The best match for the ARPES data at eV and K is found for eV (Suppl. section S2) and Å (Suppl. section S3).
The band structure is additionally unfolded to the atomic Brillouin zone (grey hexagon in Fig. 3a) Popescu and Zunger (2012) (Suppl. Fig. S2). This reflects the fact that the enlarged periodicity by AFM order is barely imprinted in the wave functions Hüfner (2013). Indeed, maps of the ARPES data did not show any patterns reminiscent of the magnetic Brillouin zone, but only of the hexagonal, geometric one (Suppl. section S1). Figure 3c shows the same DFT+U data as in b, but after selecting only the orbital contributions s, pz, and d within a frame where the -direction is pointing perpendicular to the layers. According to simplified selection rules assuming plane waves as final states Moser (2017), these orbital contributions are excited preferably for our light polarization geometry Strasdas et al. (2023). Figure 3e shows the same selection of bands for the Nèel configuration (disordered magnetic configuration in Suppl. Fig. S10), while Fig. 3d shows the ARPES curvature data below . The curvature is used for better visibility of the various features (Materials and Methods section). We always cross-check with the raw data as in Fig. 2 to corroborate that a feature is not an artifact of the curvature derivation. Orbital contributions to the different bands beyond the simplified selection rules are analyzed in Suppl. section S4 with the results included in the following discussion.
Multiple features are similar between ARPES (Fig. 3d) and DFT+U (Fig. 3c) as labeled by i-ix. The bright feature (i) and the nearly parabolic feature (ix) are used to align the Fermi level of the ARPES data to the DFT+U band structure. The additional structure in ARPES at eV is discussed below. The band (i) has a dominating S 3p character with contributions from Ni 3d (Supplementary figure S6). It is rather flat in calculation and experiment. The lowest energy band (ix) has a parabolic dispersion close to in both cases and a mixed P 3pz and S 3p character. This band appears quite similarly for MnPS3 and FePS3 Strasdas et al. (2023); Pestka et al. (2024). It is slightly lower in the DFT+U data than in the ARPES data. This is probably due to the facts that moves upwards with decreasing in ARPES at constant and that the band moves upwards with increasing in the DFT+U data (Suppl. Fig. S4/S5). At slightly larger are bands with nearly exclusive S 3px,y character (viii) that are barely visible in ARPES in accordance with the simplified selection rules Moser (2017). The remaining features ii-vii exhibit a mixed Ni 3d/S 3p character. In ARPES, they are dominated by a large cross-type structure gapped at (iv,vi) with a rather flat band on top (ii). A flat band of strong Ni 3d character (ii) and a gapped crossing of mixed character (iv,vi) are also visible in DFT+U. Other features are more subtle and partially distinct such as the relatively steep lower part of the cross from vi to vii in ARPES which is not continuous across the whole energy range in DFT+U. Nevertheless, the agreement in the whole energy range below eV is rather satisfactory regarding the fact that the band structure of the calculation was only rigidly shifted for adaption.
This importantly implies that the magnetically induced band shift identified in Fig. 2 is at the gapped crossing of feature (iv) and (vi), more precisely it is the upper part (iv) at eV around in the ARPES data (Fig. 3d) and at eV in the DFT+U data (Fig. 3c). Consistently, these band structure features exhibit a relatively strong change in DFT+U data, if the magnetic structure is changed. The gap at is increased from eV for the zigzag configuration to 0.4 eV for the Néel configuration (Fig. 3e, box). For a disordered magnetic structure, we find that features (iv,vi) get shifted upwards by meV with respect to the zigzag configuration (Suppl. Fig. S10). Data along the direction (Suppl. Fig. S9) lead to the same conclusions.
The four bands in the energy region of the observed band shift consist of all Ni 3d and S 3p orbitals as well as of a minor contribution from P 3pz (Suppl. section S4). However, only one of these bands features a bonding configuration between neighboring Ni atoms. Figure 4b shows the corresponding charge density for the state at (charge density of the other three bands, Suppl. section S5). Using the octahedral coordinate system with along a Ni-S bond, one observes Ni 3t2g orbitals (dxy, dxz, dyz) for both Ni zigzag chains, distinct by their spin orientation. On the left chain with spin , the Ni 3t2g orbitals do not have any overlap with neighboring orbitals. In contrast, for the right chain, three lobes of each Ni 3 orbital overlap with a lobe of an adjacent S 3p orbital. The zoom in Fig. 4c as well as rotated views (Suppl. movie) reveal that two of these S 3p orbitals provide a connection path between two Ni atoms (arrows in Fig. 4b), each via the two lobes of a single S 3p orbital. The third connected S 3p orbital (lower lobe of the upper Ni atom in Fig. 4c) is, however, a dead end, i. e. the S 3p orbital does not have any connection to other atoms (Suppl. movie). The final forth lobe is not connected at all and, hence, called a lone lobe as marked in Fig. 4c.
The superexchange path between 2 Ni atoms by only one S orbital on both sides implies that the interaction is antiferromagnetic. This naturally explains that the band is shifted upwards in energy when these chains become ferromagnetic within the zigzag configuration making the particular bond more unfavorable. It remains unclear why we observe only this change of a band and none of the spin-polarized 3eg levels at higher energy. We conjecture that this is related to the selection rules in our ARPES geometry not being sensitive to the corresponding levels (Suppl. Fig. S6).
We note in passing that we did not find any AFM induced C3 symmetry breaking in the band structure of NiPS3, neither by ARPES nor in the DFT+U data. The three different or directions are identical within error bars, albeit the antiferromagnetic zigzag structure breaks the C3 symmetry (Fig. 1a). Since the calculations do not reveal any symmetry breaking either, the effect is likely too weak to be detected. In ARPES, we can, moreover, not exclude that the relatively small single domains (1-10 m) Lee et al. (2024) partially overlap within the extraction spot of the photoelectrons (diameter: 5 m) He et al. (2024).
This reasonable interpretation leaves us with the question of the origin of the flat feature in ARPES at eV, that is not captured by DFT+U. It appears as a shoulder in the raw ARPES data at energies above the valence band maximum and is nearly identical above and below (Suppl. Fig S11). It, moreover, exhibits the identical photon energy () dependence as the valence band maximum (Suppl. Fig. S5, S11). To substantiate its discussion, we compare with the data of MnPS3 Strasdas et al. (2023) (Fig. 5), where all observed features in the ARPES data could be identified with bands of the DFT+U calculation. Figure 5a-f displays the direct comparison using DFT+U data within the respective antiferromagnetic ground state. The figure includes the partial density of states (PDOS) (Fig. 5a,f) where the DOS is projected to the S and P atoms (left) or the eg (3d) and t2g (3d) orbitals of Ni and Mn (right), respectively, using the octahedral coordinates with along a Ni-S bond (Fig. 5b,e). Moreover, an additional distinction between minority and majority spins within an individual atom has been performed for the Ni 3d levels. Finally, the ARPES and DFT+U data of NiPS3 are shifted upwards by 0.4 eV with respect to their to align the low energy bands of the two materials. Hence, is the Fermi level of MnPS3 only.
We find large similarities between NiPS3 and MnPS3. For example, the upper valence band of S 3p/Ni 3d character, respectively S 3p/Mn 3d character is nearly identical in strength of dispersion and orbital contributions. Moreover, the lowest energy bands ( eV) of mixed S 3p/P 3p character are very similar including the relative contribution of S and P orbitals. The energy range in between is, however, different. Firstly, there is a full spin polarization of the t2g orbitals for MnPS3 that is absent for NiPS3. Related, the mixed 3d/S 3p bands are more confined in energy for MnPS3. In particular, the two bands dominated by Mn 3d orbitals ( eV ) are the ones that are most strongly changed across Strasdas et al. (2023). In contrast, for NiPS3, all bands with eV have strong, often dominating contributions from Ni 3d orbitals (Suppl. section S4). This difference is in line with the 3d level occupation (see below). Most importantly, the experimental data is very reasonably fitted by DFT+U for both data sets except for the barely dispersing feature at eV for NiPS3 (highlighted by the yellow box in Fig. 5c). The correspondence between ARPES and DFT+U is also reasonable for FePS3, albeit a stretch of DFT+U data by 7 % was required for a quantitative match Pestka et al. (2024).
Using the octahedral coordinates along the bond directions to assign the 3d levels leads to a nice visualization of the expected 3d level occupation for the two materials (Fig. 5g-l). For Mn2+ 3d5, a complete spin polarization of the 3d levels is obtained (Fig. 5k) as expected from Hund’s rule (Fig. 5l). In addition, the occupied eg levels are split into two parts due to the deviations from a perfect octahedral symmetry Figgis and Hitchman (2000); Trimarchi et al. (2018). On the other hand, Ni2+ 3d8 shows completely occupied t2g levels, while only the eg levels are close to fully spin-polarized (Fig. 5h). Hence, the occupation of bands is in line with the expected 3d level occupations and energy splittings. Moreover, for both materials, the bands that are changing most strongly across exhibit significant 3d contributions. This makes us confident that the band assignment by DFT+U is correct and the feature at eV is beyond the related approximations. It is very likely that 3d level multiplets are involved, maybe including ligand contributions Kang et al. (2020), as will be discussed in a separate publication.
III Conclusions
We probed the band structure of NiPS3 above and below the Néel temperature using -ARPES on flakes exfoliated onto a Au surface. The ARPES data are compared with DFT+U calculations to identify the orbital character of bands. We identify an upwards band shift below by roughly 150 meV for a band close to . This band consists of mixed Ni 3d and S 3p orbitals that likely feature an antiferromagnetic Ni 3t2g-S 3p-Ni 3t2g superexchange path. Beyond the good agreement between DFT+U and ARPES that we also found for FePS3 Pestka et al. (2024) and MnPS3 Strasdas et al. (2023), we identified a shoulder in the ARPES data above the valence band maximum that could not be attributed to a band of the DFT+U data. This is in striking contrast to FePS3 and MnPS3. We assume that the additional structure is reminiscent of the known strong correlations in NiPS3.
Materials and Methods
NiPS3 crystal growth: Bulk single crystals of NiPS3 were synthesized via the vapor-transport method using stoichiometric amounts of Ni, red P, and S powders, with an additional 5 excess of sulfur to act as a transport agent. The mixture was ground to a fine powder and then sealed under high vacuum ( mbar) in a quartz ampoule. This ampoule was then placed in a two-zone furnace, set at 740 ∘C in the precursor zone and 690 ∘C in the product zone, for one week.
Sample preparation: We employ an optimized exfoliation method (scotch tape: Nitto ELP BT-150E-KL) to prepare samples with few-layered flakes of NiPS3 in a cleanroom environment. Commercially available Si/SiO2 (oxide thickness 90 nm) is used as substrate. It has been metal-coated at 370 K with Au (5 nm)/Ti (1 nm) in an ultrahigh vacuum e-beam evaporator. This improves the flake adhesion and renders the substrate conducting, which is necessary to prevent charging during ARPES Strasdas et al. (2023); Koitzsch et al. (2023); Velický et al. (2018). A mild O2 plasma ashing (power = 50 W, flux = 100 sccm) is done at room temperature for 25 s just before exfoliating NiPS3 flakes to efficiently remove organic molecules from the gold surface by oxidation. NiPS3 is exfoliated at 60∘C within few minutes after the plasma cleaning. This leads to flakes with diverse thickness, where the desired thickness and flake size are identified using commercial optical and atomic force microscopes Strasdas et al. (2023).
ARPES and XPS experiments: -ARPES and XPS were performed at the NanoESCA beamline of the Elettra synchrotron radiation facility in Italy. An ultrahigh vacuum FOCUS NanoESCA photoemission electron microscope (PEEM) was used for the -space mapping mode operation working at a background pressure of mbar Schneider et al. (2012). To ensure clean surfaces, the samples are firstly annealed at 200∘C for a couple of hours prior to ARPES. The synchrotron allows to focus a wide range of photon energies on the sample with a beam spot of 5-10 m in diameter and total energy resolution (beamline and analyzer) of 50 meV. For our experiments, the photon beam impinged with an incident angle of with respect to the surface normal and with a azimuth relative to the direction of the crystal. The polarization was within the plane of incidence (p-polarized). The Fermi level has been determined by ARPES on the Au film of the substrate. The sample temperature is adjusted by changing the liquid He flow rate and measured by a Lake Shore DT-670E-BR Si diode placed at the sample holder.
ARPES curvature: We partially use ARPES curvature plots for better visibility of the bands Zhang et al. (2011). The curvature is calculated with respect to energy after smoothing the raw data, as described in the caption of Fig. 2c, via:
and are the first and second partial derivative with respect to energy of the smoothed photoelectron intensity , is the maximum value of the first partial derivative, while is a free parameter used to adjust the relative intensity of different bands for better visibility. Here, we use .
Computational Details: The calculations were performed in the framework of density functional theory (DFT) using the generalized gradient approximation within the PBE flavor Perdew et al. (1996), as implemented in VASP software Kresse and Müller (1996). The ion–electron interactions were described by the projector augmented wave (PAW) method Holzwarth et al. (2001). Plane-wave basis cutoff and centered Monkhorst-Pack Monkhorst and Pack (1976) k-point grid were set to 550 eV and , respectively. A Gaussian smearing of 0.05 eV was employed for the Brillouin zone (BZ) integration. The interlayer vdW forces were treated within the Grimme scheme using D3 correction Grimme et al. (2010). All of the results were obtained using PBE+U method based on Dudarev’s approach Dudarev et al. (1998), with the effective on-site Hubbard U parameter (, where is fixed to eV) for the 3d orbitals. The position of the atoms and unit cell were fully optimized within the PBE+U approach. To simulate the disordered paramagnetic state, we used a supercell with spin-orbit interaction and randomly set the directions of the magnetic spin moments (in 3D) so that the net moment was zero. Here, we did not optimize the geometric structure, in order to capture only changes in electronic dispersion. Moreover, we used the widely employed band unfolding method to obtain an effective electronic structure Popescu and Zunger (2012).
Data Availability Statement
All data that support the findings of this study are available in zenodo at DOI 10.5281/zenodo.16162322.
Supporting Information
Details on sample preparation, experimental and computational methods, adaption of and including dependence of ARPES data, individual orbital projections of bands, additional charge density plots, ARPES and DFT band structure along , DFT+U band structure of paramagnetic phase and analysis of ARPES structure at eV.
Acknowledgement
We acknowledge financial support from the German Research Foundation (DFG) via the project Mo 858/19-1 and Germany’s Excellence Strategy Cluster of Excellence: Matter and Light for Quantum Computing(ML4Q) EXC 2004/1 (project Nr. 390534769) from the German Ministry of Education and Research (Project 05K2022 -ioARPES). B.B. acknowledges the support of Humboldt Research Fellowship funded by the Alexander von Humboldt Foundation. M.B. acknowledges financial support from the University of Warsaw under the "Excellence Initiative - Research University" project. PL-Grid Polish Infrastructure for Supporting Computational Science in the European Research Space, and of the Interdisciplinary Center of Modeling (ICM), University of Warsaw are gratefully acknowledged. K.W. acknowledges support by the National Science Centre in Poland under the project 2024/55/B/ST3/03144. A.K.B. and E.L. were supported by the European Commission via the Marie-Sklodowska Curie action Phonsi (H2020-MSCA-ITN-642656).